# MSCloudCAM: Multi-Scale Context Adaptation with Convolutional Cross-Attention for Multispectral Cloud Segmentation

### v3

@@ Abstract
Clouds remain a major obstacle in optical satellite imaging, limiting accurate environmental and climate analysis. To address the strong spectral variability and the large scale differences among cloud types, we propose MSCloudCAM, a novel multi-scale context adapter network with convolution based cross-attention tailored for multispectral and multi-sensor cloud segmentation. A key contribution of MSCloudCAM is the explicit modeling of multiple complementary multi-scale context extractors. And also, rather than simply stacking or concatenating their outputs, our formulation uses one extractor’s fine-resolution features and the other extractor’s global contextual representations enabling dynamic, scale-aware feature selection. Building on this idea, we design a new convolution-based cross attention adapter that effectively fuses localized, detailed information with broader multi-scale context. Integrated with a hierarchical vision backbone and refined through channel and spatial attention mechanisms, MSCloudCAM achieves strong spectral–spatial discrimination. Experiments on various multisensor datatsets e.g. CloudSEN12 (Sentinel-2) and L8Biome (Landsat-8) show that MSCloudCAM outperforms recent state-of-the-art models while maintaining competitive model complexity, highlighting the novelty and effectiveness of the proposed design for large-scale Earth observation.
@@ I Introduction
Figure 1: High-level architecture of MSCloudCAM. The model integrates a hierarchical vision backbone with dual multi-scale context extractors and introduces a novel cross-attention module.
To address these challenges, we propose MSCloudCAM, a cross-attention with multi-scale context network designed for robust cloud segmentation in multispectral, multi-sensor imagery. MSCloudCAM leverages complementary multi-scale context encoders and a newly designed cross-attention mechanism to dynamically fuse fine-grained spatial features with global contextual information. This design enables accurate pixel-wise classification across datasets with differing spectral responses, resolutions, and cloud characteristics.
Our key contributions are as follows:
@@
  • We design a novel convolution-based cross-attention module, i.e. CAM that leverages these complementary feature hierarchies through an adaptive formulation. This mechanism enables dynamic, scale-aware feature fusion, offering a more principled alternative to conventional stacking or concatenation.
@@ II-B Deep Learning Approaches
Deep learning has become the dominant strategy for cloud segmentation due to its ability to model spatial–spectral structure directly from multispectral inputs.
CNN-based architectures such as UNet [6] and UNetMobv2 [7] capture fine localization through skip connections.DeepLabV3+ [8] enhances multi-scale context via dilated convolutions and ASPP. Numerous CNN variants target cloud detection specifically, including CDNetv1/v2 [9, 10], KappaMask [11], DBNet [12], SCNN [13], MCDNet [14], and HRCloudNet [15], each improving spectral–spatial fusion or multi-scale representation.
@@ II-C Gaps in Current Methods
Despite significant progress, most approaches are designed for single-sensor datasets, limiting cross-sensor generalization. Moreover, few models explicitly integrate complementary multi-scale spectral–spatial cues using directional cross-attention. Our MSCloudCAM addresses both challenges.
@@ III-A Overview
MSCloudCAM is a multispectral semantic segmentation network that integrates a hierarchical vision backbone e.g. Swin Transformer [19] encoder with dual context extractor modules, a cross-attention fusion mechanism, and a lightweight combined attention refinement block. The model is designed to capture long-range spectral–spatial dependencies while retaining fine-grained local structure. A multi-stage decoder with auxiliary supervision produces the final dense predictions. The overall pipeline is illustrated in Fig. 2.
Figure 2: Architecture of MSCloudCAM. A multispectral Swin Transformer encoder generates hierarchical features. ASPP and PSP modules enrich deep and intermediate representations. These are fused by convolutional cross-attention, refined via channel–spatial attention, and decoded by a multi-stage upsampling path with auxiliary supervision.
@@ III-B Backbone Feature Extraction
Let $\mathbf{X}\in\mathbb{R}^{B\times C_{\text{in}}\times H_{0}\times W_{0}}$ denote the input multispectral image with $C_{\text{in}}$ spectral channels (13 for Sentinel-2, 11 for Landsat-8). The encoder $\mathcal{E}(\cdot)$ produces a hierarchy of features:
@@
with channel dimensions $[96,192,384,768]$ and progressively reduced spatial sizes $(H_{i},W_{i})$. Shifted-window self-attention within Swin ensures that $f_{4}$ captures global dependencies, while $f_{1}$$f_{3}$ preserve finer detail.
@@ III-C Multi-Scale Context Modules
As a part of dual multiscale context extractor, ASPP [8] is applied to the deepest feature map $f_{4}$, Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
@@
where $\oplus$ denotes channel-wise concatenation, $r$ specifies the dilation rate, GAP is global average pooling, and $\sigma$ represents the ReLU nonlinearity. This module injects large-scale semantic context while retaining spatial discrimination.
For the intermediate feature map $f_{3}$, Pyramid Scene Parsing i.e. PSP [21] aggregates multi-scale contextual cues by adaptive pooling at several grid scales:
@@
where $\text{Pool}_{s}$ denotes average pooling to an $s\times s$ grid, followed by projection and bilinear upsampling. This captures scene-level information at progressively coarser scales, which is essential for delineating fine structures such as thin clouds.
III-D Cross-Attention Fusion
@@
where $x_{\text{MC1}}$ emphasizes fine spatial detail from ASPP module and $x_{\text{MC2}}$ from PSP module encodes broader contextual semantics. To effectively combine these complementary cues, we employ a convolutional cross-attention mechanism. Here, $x_{\text{cat}}$ is used to generate the queries, as the concatenated representation benefits from selectively incorporating global context to refine local structure. Meanwhile, $x_{\text{MC2}}$ provides the keys and values, since its pooled multi-scale features offer stable, region-level guidance suitable for modulating finer details. This directional interaction allows each spatial location in $x_{\text{MC1}}$ to attend to the most relevant semantic context encoded by $x_{\text{MC2}}$.
@@ III-G Decoder with Deep Supervision
IV Experiments
TABLE I: Performance comparison on CloudSEN12 L2A (BOA).
@@ IV-A Datasets
CloudSEN12 [1] is a global Sentinel-2 cloud segmentation dataset with 13 spectral bands (10–60 m) and pixel-level labels for clear (0), thick cloud (1), thin cloud (2), and cloud shadow (3). We use the high subset, selecting 10k samples each from L1C and L2A. L1C provides Top-of-Atmosphere (TOA) reflectance, while L2A provides Bottom-of-Atmosphere (BOA) surface reflectance. All bands are normalized by dividing top-of-atmosphere reflectance by $3000.0$.
L8Biome [2] is derived from Landsat-8 CCA, containing 11 bands (30 m) with manual masks. The original five classes—fill (0), shadow (64), clear (128), thin cloud (192), and cloud (255)—are remapped to CloudSEN12’s taxonomy: clear $\rightarrow 0$, thick cloud $\rightarrow 1$, thin cloud $\rightarrow 2$, shadow $\rightarrow 3$. The fill class is ignored. Bands are normalized with the same $3000.0$ divisor for consistency.
TABLE II: Dataset split statistics for CloudSEN12 and L8Biome.
@@ IV-D Evaluation Metrics
Class-wise metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes, giving equal weight regardless of class size. In contrast, overall accuracy (aAcc) is computed over all pixels, thus weighting larger classes more heavily.
V Results
V-A Quantitative Performance
We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables IIIIV present the detailed results.
TABLE III: Performance comparison on CloudSEN12 L1C (TOA).
Model IoU (%) F1 (%) Accuracy (%) aAcc
U-Net-MobV2 90.82 85.98 50.54 67.21 73.64 95.19 92.46 67.14 80.39 83.80 94.42 96.11 95.05 96.62 95.55 91.10
DeepLabV3+-ResNet101 90.65 85.66 50.60 66.36 73.32 95.10 92.28 67.20 79.78 83.59 94.38 95.94 94.94 96.47 95.43 90.87
@@ V-A Quantitative Performance
V-B Qualitative Results
To further demonstrate the effectiveness of our method, Fig. 3 presents visual comparisons of cloud segmentation results. Each row shows a different dataset, and each column a different method.
Figure 3: Qualitative cloud segmentation results. Each row corresponds to a dataset. The proposed method produces sharper delineation of thin clouds and shadows compared to other approaches. Class–color mapping: Class 1 (Clear sky – light blue), Class 2 (Thick cloud – white), Class 3 (Thin cloud – light gray), Class 4 (Cloud shadow – dark gray).
V-C Ablation Study
@@
Using PSP and ASPP together provides complementary global–local context and outperforms either module alone. Removing the Cross-Attention Module and using simple concatenation weakens the model’s ability to emphasize informative regions. The full MSCloudCAM system with PSP, ASPP, and CAM delivers the highest accuracy.
TABLE V: Ablation study evaluating dual multi context extractor configurations and our cross-attention module (CAM) on CloudSEN12 L2A (BOA).
Dual Extractor Fusion Type mIoU mDice mAcc aAcc
PSP + PSP CAM 70.26 81.61 92.05 86.41
ASPP + ASPP CAM 70.09 81.25 91.93 86.18
PSP + ASPP Concat 71.48 82.44 93.52 88.02
PSP + ASPP CAM 73.12 83.70 94.85 89.69
V-D Model Complexity Analysis
Table VI summarizes the computational complexity of all evaluated models in terms of FLOPs, parameter counts, and inference efficiency. In addition to latency (ms/image), we report throughput (TP, samples/s) to quantify batch inference performance. All experiments were conducted on a single NVIDIA A100 40GB GPU using PyTorch 2.3 and CUDA 12.1, with a batch size of 8 and an input resolution of 512×512. All models were evaluated in FP32 precision. Our proposed MSCloudCAM achieves a noteworthy balance between accuracy and computational efficiency and CAM module only adds 1.7M trainable parameters to the model.
@@ V-D Model Complexity Analysis
Mask2Former-DINOv2-B14 85.6 90.6 44.0 22.7
@@
VI Code and Model Availability
The full implementation of our proposed MS-CloudCAM framework is publicly available. The source code, training scripts, and experimental configurations are hosted on GitHub: https://github.com/mazid-rafee/ms-cloudcam. Additionally, pretrained model checkpoints are released on Hugging Face: https://huggingface.co/mazid-rafee/MS-CloudCAM/tree/main.
VII Conclusion
We presented MSCloudCAM, a cross attention with multi-scale context network for multispectral, multi-sensor cloud segmentation. By integrating hierarchical vision backbone with dual multi-scale context extractors and emlpoying our novel convolution based cross attention module, our method achieves superior segmentation accuracy over state-of-art-models on CloudSEN12 and L8Biome datasets. Future work will explore lightweight variants for onboard satellite processing and extend the model to spatiotemporal cloud tracking.

### v4

@@ Abstract
Clouds remain a major obstacle in optical satellite imaging, limiting accurate environmental and climate analysis. To address the strong spectral variability and the large scale differences among cloud types, we propose MSCloudCAM, a novel multi-scale context adapter network with convolution based cross-attention tailored for multispectral and multi-sensor cloud segmentation. A key contribution of MSCloudCAM is the explicit modeling of multiple complementary multi-scale context extractors. Our formulation uses one extractor’s fine-resolution features and the other extractor’s global contextual representations enabling dynamic, scale-aware feature selection. Building on this idea, we design a new convolution-based cross attention adaptation module (CAM) that effectively fuses localized, detailed information with broader multi-scale context rather than simply stacking or concatenating context extractors’ outputs. Integrated with a hierarchical vision backbone and refined through channel and spatial attention mechanisms, MSCloudCAM achieves strong spectral–spatial discrimination. Experiments on multisensor datasets e.g. CloudSEN12 (Sentinel-2) and L8Biome (Landsat-8), demonstrate that MSCloudCAM achieves superior overall segmentation performance and competitive class-wise accuracy compared to recent state-of-the-art models, while maintaining efficient model complexity, highlighting the effectiveness of the proposed design for large-scale Earth observation.
Index Terms—Cloud segmentation, Multispectral Imagery, Deep Learning, Convolutional Cross Attention, Multi-scale Context, Sentinel-2, Landsat-8.
@@ I Introduction
Figure 1: High-level architecture of MSCloudCAM. The model integrates a hierarchical vision backbone with dual multi-scale context extractors and introduces a novel convolutional cross-attention adaptation module (CAM).
To address these challenges, we propose MSCloudCAM, a cross-attention adapter with multi-scale context network designed for robust cloud segmentation in multispectral, multi-sensor imagery. Our key contributions are as follows:
@@
  • We design a novel convolution-based cross-attention adaptation module, i.e. CAM that leverages these complementary feature hierarchies through an adaptive formulation. This mechanism enables dynamic, scale-aware feature fusion, offering a more principled alternative to conventional stacking or concatenation.
@@ II-B Deep Learning Approaches
CNN-based deep learning architectures such as UNet [6] and UNetMobv2 [7] capture fine localization through skip connections.DeepLabV3+ [8] enhances multi-scale context via dilated convolutions and ASPP. Numerous CNN variants target cloud detection specifically, including CDNetv1/v2 [9, 10], KappaMask [11], DBNet [12], SCNN [13], MCDNet [14], and HRCloudNet [15], each improving spectral–spatial fusion or multi-scale representation.
@@ II-C Gaps in Current Methods
Despite significant progress, most approaches are designed for single-sensor datasets, limiting cross-sensor generalization. Moreover, few models explicitly integrate complementary multi-scale spectral–spatial cues using directional cross-attention adaptation. Our MSCloudCAM addresses both challenges.
@@ III-A Overview
MSCloudCAM is a multispectral semantic segmentation network that integrates a hierarchical vision backbone e.g. Swin Transformer [19] encoder with dual context extractor modules, a cross-attention adaptation mechanism, and a lightweight (only adds 1.7M trainable parameters) combined attention refinement block. The model is designed to capture long-range spectral–spatial dependencies while retaining fine-grained local structure. A multi-stage decoder with auxiliary supervision produces the final dense predictions. The overall pipeline is illustrated in Fig. 2.
Figure 2: Architecture of MSCloudCAM. A multispectral Swin Transformer encoder generates hierarchical features. ASPP and PSP modules enrich deep and intermediate representations. These are fused by convolutional cross-attention adapter, refined via channel–spatial attention, and decoded by a multi-stage upsampling path with auxiliary supervision.
@@ III-B Backbone Feature Extraction
Let $\mathbf{X}\in\mathbb{R}^{B\times C_{\text{in}}\times H_{0}\times W_{0}}$ denote the input multispectral image with $C_{\text{in}}$ spectral channels (13 for Sentinel-2, 11 for Landsat-8), where $B$ is the batch size and $(H_{0},W_{0})$ denote the input spatial resolution. The encoder $\mathcal{E}(\cdot)$ produces a hierarchy of features:
@@
where $(H_{i},W_{i})$ denote the spatial resolutions at stage $i$, progressively reduced from $(H_{0},W_{0})$ through hierarchical downsampling, and $C_{i}\in\{96,192,384,768\}$. Shifted-window self-attention within Swin ensures that $f_{4}$ captures global dependencies, while $f_{1}$$f_{3}$ preserve finer spatial details.
@@ III-C Multi-Scale Context Modules
As part of the dual multiscale context extractor, ASPP [8] is applied to the deepest feature map $f_{4}$. Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
@@
where $\oplus$ denotes channel-wise concatenation, $r$ specifies the dilation rate, GAP is global average pooling, and $\sigma$ represents the ReLU nonlinearity. This operation aggregates contextual information across multiple receptive fields.
For the intermediate feature map $f_{3}$, Pyramid Scene Parsing (PSP) [21] aggregates contextual cues by adaptive pooling at several grid scales:
@@
where where $s\in\{1,2,3,6\}$ denotes the pooling grid size and $\text{Pool}_{s}$ maps $f_{3}$ to an $s\times s$ grid via average pooling, followed by projection and bilinear upsampling. This operation encodes multi-scale scene layout information.
This asymmetric placement aligns contextual operators with the feature hierarchy. PSP is applied to $f_{3}$ to exploit its higher spatial resolution for preserving structural details and accurate boundary localization, whereas ASPP is applied to the deeper $f_{4}$ where stronger semantic abstractions and a coarser resolution benefit more from large-dilation convolutions and global context aggregation.
III-D Cross-Attention Adaptation
@@
where $x_{\text{MC1}}$ emphasizes fine spatial detail from ASPP module and $x_{\text{MC2}}$ from PSP module encodes broader contextual semantics. To effectively combine these complementary cues, we employ a convolutional cross-attention adaptation mechanism. Here, $x_{\text{cat}}$ is used to generate the queries, as the concatenated representation benefits from selectively incorporating global context to refine local structure. Meanwhile, $x_{\text{MC2}}$ provides the keys and values, since its pooled multi-scale features offer stable, region-level guidance suitable for modulating finer details. This directional interaction allows each spatial location in $x_{\text{MC1}}$ to attend to the most relevant semantic context encoded by $x_{\text{MC2}}$.
@@ III-G Decoder with Deep Supervision
TABLE I: Performance comparison on CloudSEN12 L1C (TOA).
Model IoU (%) F1 (%) Accuracy (%) aAcc
UNet-MobV2 90.82 85.98 50.54 67.21 73.64 95.19 92.46 67.14 80.39 83.80 94.42 96.11 95.05 96.62 95.55 91.10
TABLE II: Performance comparison on CloudSEN12 L2A (BOA).
@@ III-H Datasets
CloudSEN12 [1] is a global Sentinel-2 cloud segmentation dataset with 13 spectral bands (10–60 m) and pixel-level labels for clear (0), thick cloud (1), thin cloud (2), and cloud shadow (3). We use the high subset, selecting 10k samples each from L1C (Top-of-Atmosphere - TOA) and L2A ( Bottom-of-Atmosphere - BOA). All bands are normalized by dividing top-of-atmosphere reflectance by $3000.0$.
L8Biome [2] is derived from Landsat-8 Cloud Cover Assessment, containing 11 bands (30 m) with manual masks. The original five classes—fill (0), shadow (64), clear (128), thin cloud (192), and cloud (255)—are remapped to CloudSEN12’s taxonomy: clear $\rightarrow 0$, thick cloud $\rightarrow 1$, thin cloud $\rightarrow 2$, shadow $\rightarrow 3$. The fill class is ignored. Bands are normalized with the same $3000.0$ divisor for consistency.
TABLE III: Dataset split statistics for CloudSEN12 and L8Biome.
@@ III-K Evaluation Metrics
Here, $\mathrm{TP}_{i}$, $\mathrm{FP}_{i}$, $\mathrm{FN}_{i}$, and $\mathrm{TN}_{i}$ denote the numbers of true positives, false positives, false negatives, and true negatives for class $i$, respectively. Class-wise metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes. In contrast, overall accuracy (aAcc) is computed over all pixels.
IV Results
IV-A Quantitative Performance
We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables I, II, IV present the detailed results.
@@ IV-A Quantitative Performance
IV-B Qualitative Results
To further demonstrate the effectiveness of our method, Fig. 3 presents visual comparisons of cloud segmentation results.
Figure 3: Qualitative cloud segmentation results across datasets. Each row represents a dataset.The proposed method shows sharper delineation of thin clouds and shadows than other approaches. Class–color mapping: Class 1 (Clear sky – light blue), Class 2 (Thick cloud – white), Class 3 (Thin cloud – light gray), Class 4 (Cloud shadow – dark gray)
IV-C Ablation Study
@@
Using PSP and ASPP together provides complementary global–local context and outperforms either module alone. Removing the cross-attention module (CAM) and using simple concatenation weakens the model’s ability to emphasize informative regions. The full MSCloudCAM system with PSP, ASPP, and CAM delivers the highest accuracy. Overall, these findings confirm that each component contributes uniquely to the system.
TABLE V: Ablation study evaluating dual multi context extractor and cross-attention module (CAM) on CloudSEN12 L1C, CloudSEN L2A, L8Biome datasets.
Dataset Extractor Fusion mIoU mF1 mAcc
CloudSEN12 PSP + PSP CAM 72.31 83.56 94.11
L1C ASPP + ASPP CAM 71.98 83.37 93.89
PSP + ASPP Concat 73.43 84.29 94.75
PSP + ASPP CAM 75.52 85.13 95.99
CloudSEN12 PSP + PSP CAM 70.26 81.61 92.05
L2A ASPP + ASPP CAM 70.09 81.25 91.92
PSP + ASPP Concat 71.48 82.44 92.63
PSP + ASPP CAM 73.12 83.70 94.85
L8Biome PSP + PSP CAM 54.95 67.63 91.87
ASPP + ASPP CAM 55.08 67.31 91.46
PSP + ASPP Concat 56.41 68.22 92.02
PSP + ASPP CAM 58.04 69.18 93.55
IV-D Model Complexity Analysis
Table VI summarizes the computational complexity of evaluated models in terms of FLOPs, parameter counts, and inference efficiency. In addition to latency (ms/image), we report throughput (TP, samples/s) to quantify batch inference performance. All experiments were conducted on a single NVIDIA A100 40GB GPU using PyTorch 2.3 and CUDA 12.1, with a batch size of 8 and an input resolution of 512×512. All models were evaluated in FP32 precision. Our proposed MSCloudCAM achieves a noteworthy balance between accuracy and computational efficiency and CAM only adds 1.7M trainable parameters to the model.
@@ IV-D Model Complexity Analysis
Mask2Former-DINOv2 85.6 90.6 44.0 22.7
@@
V Code and Model Availability
The full implementation of our proposed MSCloudCAM framework is publicly available. The source code, training scripts, and experimental configurations are hosted on GitHub: https://github.com/mazid-rafee/ms-cloudcam. Additionally, pretrained model checkpoints are released on Hugging Face: https://huggingface.co/mazid-rafee/MS-CloudCAM/tree/main.
VI Conclusion
We presented MSCloudCAM which integrates hierarchical vision backbone with dual multi-scale context extractors and employs our novel convolution based cross attention adaptation module. Extensive experiments on the CloudSEN12 and L8Biome datasets demonstrate that MSCloudCAM consistently surpasses state-of-the-art methods across multiple segmentation metrics. These highlight the effectiveness and generalizability of our design for multi-sensor and multi-spectral cloud understanding. While our experiments focus on transformer-based backbones, evaluating the proposed context adapter with CNN-based encoders remains an interesting direction for future work.